OpenStudy (anonymous):

I need help finding an equation of the line that is parallel to the line 2x-3y=-4; containing the point (5,-3) please!!

OpenStudy (skullpatrol):

Find the slope of 2x-3y=-4 by putting it into the y=mx+b form. y=(2/3)x +(4/3) Now you have the slope of 2/3 and a point (5,-3). Can you find the new equation?

OpenStudy (anonymous):

no.. the answer is \[y=\frac{ 2 }{3 }x +\frac{ 19 }{ 3 }\] but i don't understand where 19 came from

OpenStudy (skullpatrol):

y=mx+b -3=(2/3)(5)+b solve for b, -3-(10/3)=b -19/3=b

OpenStudy (skullpatrol):

y=(2/3)x-(19/3)